OpenStudy (anonymous):

Find the Slope intercept form. (3,-6) and (1,5)

OpenStudy (anonymous):

First find the slope of those two points

OpenStudy (anonymous):

yes, find the slope using the slope formula, them plug it into slope intercept form: y=mx+b ~~>M being the slope

OpenStudy (anonymous):

Is it -11/2

OpenStudy (anonymous):

Yep. now do point slope form: y-y1=m(x-x1)

OpenStudy (anonymous):

(x1, y1) is either one of the points you choose and m is slope

OpenStudy (anonymous):

can you do that for me I dont understand

OpenStudy (anonymous):

Lets use the point (1, 5) where 1 is x1 and 5 is y1 and plug it into the formula: y-5=-11/2(x-1)

OpenStudy (anonymous):

But the slope is just -11/2

OpenStudy (anonymous):

Yes i plugged that in for m. were trying to get slope intercept form

OpenStudy (anonymous):

Now distribute the -11/2 into the parentheses

OpenStudy (anonymous):

which will give you: y-5=-11/2x+11/2 then add the 5 to both sides and you get: y=-11/2x+21/2